# Improving Reward Models with Synthetic Critiques

### v1

@@ 6 Conclusion
We propose an accessible and cost-effective approach to improve reward models (RMs) in RLHF by enriching them with synthetic natural language critiques. We prompt LLMs to generate critiques that assess the prompt-completion pair and train RMs to predict scalar rewards conditioned on these critiques. Experiments demonstrate the effectiveness of using critiques for RMs, as evidenced by higher scores on existing RM benchmarks. We found that the benefit of critiques is greater when they are of high quality, especially for RMs initialized from weaker pretrained models and in low-data settings. In these settings one high-quality critique-enhanced preference pair is worth 40 non-enhanced preference pairs. As these critiques are generated without human labour, this approach could make it much more cost-effective to obtain competitive reward models. In our experiments, performance was still better with 100k examples, and both approaches had near equivalent performance only at 200k examples. In future work, we plan to investigate generating critiques using chain-of-thought (CoT) prompting and investigating their potential to improve LLM reasoning abilities.
@@ 7 Limitations
While we evaluate RMs using ranking test accuracy, we did not experiment with optimizing LLMs with such critique-enriched RMs, but focused on evaluating RMs directly on existing RM benchmarks. Regarding critique quality, the critiques generated by the LLMs might be incorrect, or contain hallucinations. As mentioned in § 5.2, the added value of critiques diminishes in the abundant data regime. However, even in that scenario adding critiques does not reduce the final performance and obtains better score in the Hard setting, indicating better generalization. A final limitation is the limited improvement when starting from a strong preference tuned model, which is similar to the case of abundant preference data. We would also like to extend our experiments to include a wider pretrained checkpoints, and prompting methods to understand their effects on critiques.
@@ Appendix A Prompt Template
We show the template we use to prompt LLMs for point-wise critiques in Figure A.
@@ Appendix D Scaling Performance By Task
We illustrate a break down of RewardBench Score v.s. training examples by task type in Figure 6(a). We observe that RMs enriched with Command R+ Critiques consistently outperforms No-Critiques on Reasoning task. Chat Hard contains challenging prompt completions that are designed with superficial features to trick LLM-based evaluators. On this dataset, training with Command R+ Critiques leads to a less performance drop compared to No-Critiques over increasing number of training examples, suggesting that critiques help with robustness against adversarial examples.

### v2

@@ 6 Conclusion
We propose an accessible and cost-effective approach to improve reward models (RMs) in RLHF by enriching them with synthetic natural language critiques. We prompt LLMs to generate critiques that assess the prompt-completion pair and train RMs to predict scalar rewards conditioned on these critiques. Experiments demonstrate the effectiveness of using critiques for RMs, as evidenced by their stronger performance on existing RM benchmarks. We found that critiques are especially effective in low-data settings, when they are of high quality, the base RM is weak, and the task requires reasoning skills. In these cases, a single high-quality critique-enhanced preference pair can be as valuable as 40 standard preference pairs. As these critiques are generated without human labour, this approach offers a more cost-effective way to obtain competitive reward models, given the current scarcity and high cost of preference training data. We also show that our method of leveraging critiques as part of the reward model’s Chain-of-Thought enhances the training process and interpretability. In future work, we plan to investigate their potential to improve LLM reasoning abilities.
@@ 7 Limitations
While we evaluate RMs using ranking test accuracy, we did not experiment with optimizing LLMs with such critique-enriched RMs, but focused on evaluating RMs directly on existing RM benchmarks. However, previous work (Shen et al., 2023a) suggests that stronger RMs lead to enhanced performance of LLMs during RLHF. Regarding critique quality, the critiques generated by the LLMs might be incorrect, or contain hallucinations. As mentioned in § 5.2, the added value of critiques diminishes in the abundant data regime. However, even in that scenario adding critiques does not reduce the final performance and obtains better score in the more challenging settings such as Reasoning and Chat Hard, indicating better generalization. A final limitation is the limited improvement when starting from a strong preference tuned model, which is similar to the case of abundant preference data. We would also like to extend our experiments to include a wider set of pretrained checkpoints, and prompting methods to understand their effects on critiques.
@@ Appendix A Prompt Template
We show the template we use to prompt LLMs for point-wise critiques in Fig. A. We made minimal changes starting from the template in Cui et al. (2023).
@@ Appendix D Scaling Performance By Task
We illustrate a break down of RewardBench Score v.s. training examples by task type in Fig. 6(a). We observe that RMs enriched with Command R+ Critiques consistently outperforms No-Critiques on Reasoning task. Chat Hard contains challenging prompt completions that are designed with superficial features to trick LLM-based evaluators. On this dataset, training with Command R+ Critiques leads to a less performance drop compared to No-Critiques over increasing number of training examples, suggesting that critiques help with robustness against adversarial examples. We observe that the classifier-based RMs overwhelmingly outperform the generative judge baseline. While the generative judge performs strongly on Chat, comparably on Safety, it performs poorly in Reasoning, followed by Chat Hard. Upon closer inspection on Reasoning, we see that many of the mistakes for Gen-Command R+ (score of 0.060.06~{}0.060.06) are due to the generative judge assigning the same score to both preferred and worse, leading to a tie. We hypothesize that this is because without preference training, completion pairs from the more difficult Reasoning subset appear similar to the pretrained checkpoint, which is not sensitive to their differences. However, when conditioning on critiques, Gen-Command R+ (+Critiques) spreads out the reward scores on those ties, leading to a $0.3$ increase (score of $0.3685$).